A farmer has determined that a crop of strwberries yields a yearly profit of $1.50 per square yard. If strawberries are planted on a triangul;ar piece of land whose sides are 50 yards, 75 yards, and 100 yards, how much profit to the nearest hundred dollars, would the farmer expect to make from this piece of land during the next harvest.
Answer by checkley75(3666)

because this is not a right triangle we need to use the following formula to find the area:
A=SQRT[S(S-A)(S-B)(S-C)] WHERE S=1/2 THE PERIMETER AND A,B,C ARE THE THREE SIDES.
S=(50+75+100)/2=225/2=112.5
A=SQRT[112.5(112.5-50)(112.5-75)(112.5-100)]
A=SQRT[112.5*62.5*37.5*12.5]
A=SQRT(3,295,898.3)
A=1,815.46 SQUARE YARDS OF GROUND.
THUS 1,815.46*1.50=$2,723.19 OR $2,700 PROFIT.